OpenStudy (anonymous):

Suppose f(x) is continuous and differentiable. Must f'(x) be differentiable? Explain why or give an example why not.

OpenStudy (freckles):

Trying to think of an example of when it is false. Try differentiating this function f twice: f(x)=x^2 if x>=0 and -x^2 if x<0

OpenStudy (freckles):

@Satyaprakash2009rta

OpenStudy (perl):

the answer is no

OpenStudy (perl):

try f(x) = x^(4/3) this is differentiable for all x (and therefore continuous) but f ' (x) = 4/3 x^(1/3) is not differentiable for x = 0

OpenStudy (perl):

@freckles can you help with his older question prove that lim f(c*x) = lim f(x) as x -> 0 for both limits